Your car and the trailer it is towing are initially at rest. You apply an acceleration of 1.70 m/s2 to get the vehicles moving. The mass of your car is 1250 kg and the mass of the trailer is 540 kg. (For the following, enter the magnitudes.)


 (a) Determine the net force (in N) acting on the car. N

 (b) How much force (in N) does the trailer experience? N 

 (c) Determine the force (in N) the trailer exerts on the car. N
Answer #1

Force = mass * accelaration

so mass of cart is 1250 kg ,first we will see forces on it and the find net force

X-axis force is

F=ma

Y-axis forces is

N-mg=0 ,

so N= mg we can say that for cart thats force is cancel out

If they are accelerating at 1.70 m/s^2 then the net force on them is given by F=ma, everything else cancels out (force of gravity on the car, reactive force of the road surface etc.)

The force exerted by the trailer on the car is the same as that exerted by the car on the trailer, which is the F=ma for the trailer above.

Now force on cart is

(a)F=ma=1250*1.70=2125 N

(b))F=ma=540*1.70=918 N

(c)F=ma=540*1.70=918 N

N